5(r+4)=d
12(r-4)=d
r is rate, and that is what I am trying to solve for.

Answer by mananth(16946)

5(r+4)=d
12(r-4)=d


Therefore
5(r+4)=12(r-4)
5r+20=12r-48
7r=68
r=68/7